# Graph Neural Networks are Heuristics

### v1

@@ 2 Background
3 Our Contribution
We show that encoding the Hamiltonian cycle constraint of the TSP directly into the learning objective is sufficient to achieve competitive performance without reliance on search. Our method is purely unsupervised, requiring no ground-truth solutions or reinforcement rewards.
To achieve this, we introduce three elements. First, we employ an equivariant and invariant representation that preserves permutation symmetry while removing nuisance variations such as translation and rotation, ensuring that the model operates directly on the intrinsic structure of the problem. Second, we use dropout as a source of controlled stochasticity within the network, which improves robustness without introducing iterative refinement. Third, we adopt snapshot ensembling along a single training trajectory, allowing multiple functionally distinct solutions to be obtained without additional training cost.
@@ 5.1 Equivariant Coordinate Feature Extraction
We introduce an equivariant coordinate feature extractor that maps a set of planar points to per-point features while respecting the symmetries of the Euclidean group. The construction is fully deterministic, permutation-equivariant, and (almost everywhere) invariant to global translations and rotations. The equivariant feature extraction procedure underlying these properties is illustrated in Figure 1.
5.1.1 Problem Setting
@@
5.1.2 Intrinsic Coordinates and Harmonic Features
@@
5.1.3 Equivariance and Invariance Properties
@@
$f_{\mathrm{GNN}}(T_{t}\cdot X,A(T_{t}\cdot X))=f_{\mathrm{GNN}}(X,A(X)),$ (20)
@@
$X=\{x_{i}\}_{i=1}^{n},\quad x_{i}\in\mathbb{R}^{2}$ Canonicalization $c=\frac{1}{n}\sum_{i}x_{i},\;\tilde{x}_{i}=x_{i}-c;\;\Sigma=\frac{1}{n}\sum_{i}\tilde{x}_{i}\tilde{x}_{i}^{\top};\;(u_{\perp},u)=\mathrm{eig}(\Sigma)$ (sign-fixed) Intrinsic coordinates $a_{x}(i)=\tilde{x}_{i}^{\top}u,\;\;a_{y}(i)=\tilde{x}_{i}^{\top}u_{\perp};\;\;r(i)=\sqrt{a_{x}(i)^{2}+a_{y}(i)^{2}+\varepsilon},\;\;\theta(i)=\mathrm{atan2}(a_{y}(i),a_{x}(i))$ Angular harmonics $\begin{aligned} \phi_{m}^{\sin}(i)&=\sin(m\theta(i)),\\ \phi_{m}^{\cos}(i)&=\cos(m\theta(i)),\\ m&=1,\dots,M\end{aligned}$ $f(i)=[\,r(i),a_{x}(i),a_{y}(i),\{\phi_{m}^{\sin}(i)\}_{m=1}^{M},\{\phi_{m}^{\cos}(i)\}_{m=1}^{M}\,]$ $\begin{aligned} F(X)=\{f(i)\}_{i=1}^{n}\end{aligned}$
Figure 1: Illustration of the coordinate feature extractor: a canonical frame yields intrinsic polar coordinates and Fourier harmonics, concatenated into per-point features.
@@ 6 Training and Inference
We train our scattering-attention GNN using distributed data parallelism on multi-node GPU clusters. For TSP100 and TSP200, we launch runs on $2$ nodes with $4$ GPUs per node (8 GPUs total), using per-GPU batch size $256$ and the Adam optimizer with weight decay $2.5\times 10^{-5}$, dropout $0.1$, and a learning rate $8\times 10^{-3}$. We train for up to $1000$ epochs for TSP 100 and 200, and $500$ epochs for TSP 500, with a scheduler and a $15$-epoch warmup, enable adaptive gradient clipping, and apply early stopping with patience $100$; checkpoints are saved every $5$ epochs with automatic resume enabled. For TSP500, we scale to $4$ nodes with $4$ GPUs per node (16 GPUs total) and train a larger model (hidden dimension $384$, $48$ layers) with dropout $0.5$, learning rate $2\times 10^{-3}$, weight decay $10^{-4}$, and temperature $\tau=3.5$, using a smaller per-GPU batch size of $50$ due to memory constraints. Across all settings, training and validation instances are sampled from uniform Euclidean TSP distributions, and we fix the distance scale to $5.0$ to match the inference-time construction of the graph adjacency.
@@ 7.1 Length Distribution
Figures 3 show the tour-length distributions on the test set for 100-, 200-, and 500-node instances, using the model that achieves the lowest validation length across all hyperparameter settings. Across all problem sizes, our method consistently outperforms the Greedy Nearest Neighbor (NN) baseline, which constructs tours by iteratively selecting the closest unvisited node. The resulting histograms show that our model produces shorter and more tightly concentrated tour-length distributions, with mean lengths of $\mu=8.18$, $11.65$, and $18.84$, compared to Greedy NN’s $\mu=9.67$, $13.43$, and $20.76$, respectively.
@@
500 Mean length 18.67 18.72 18.72 18.73 18.73 18.72 18.73 18.73 18.72 18.72 18.54
Wins (%) 16.3 8.8 8.8 9.2 9.5 9.7 10.0 8.5 9.7 9.5
@@
500 Mean length 18.67 18.76 18.72 18.72 18.70 18.52
Wins (%) 29.3 13.6 17.3 17.5 22.3
@@
GNN, UL 8.11 4.65% 3.37ms 11.57 8.43% 9.48ms 18.54 12.57% 59.34ms
Table 2 illustrates the behavior of our snapshot ensemble across problem sizes. Each column corresponds to a single inference run using the same model architecture, but taken from different training epochs. For TSP-100 and TSP-200, snapshots are drawn from training epochs $\{850,900,950,1000\}$, while for TSP-500 they are taken from $\{350,400,450,500\}$.
@@
Compared to other heuristics, the proposed method exhibits a consistently more favorable trade-off between solution quality and inference cost As reported in Table 3, classical heuristics such as Christofides achieve optimality gaps of 12.12%, 12.75%, and 13.30% on TSP 100, 200 and 500, respectively. In contrast, our purely feed-forward GNN heuristic attains substantially lower gaps of 5.55% and 9.18% on TSP 100 and 200 in the deterministic setting, and further reduces the gap to 4.77% and 8.83% using snapshot ensembling. On TSP 500, our snapshot and MC-dropout ensembles achieve gaps of 12.47% and 12.57%, closely matching Christofides while operating at millisecond-level inference time on GPUs. Heuristics such as 2-opt reduce this gap, but at the cost of substantially higher runtime, particularly on larger instances. These results demonstrate that learned, structure-aware heuristics can match or exceed the performance of carefully engineered classical methods, while maintaining dramatically lower computational overhead.

### v2

@@ 3.3 Hamiltonian Cycle Ensemble
4 Our Contribution
In this paper, we propose a single-model, unsupervised framework that encodes the Hamiltonian cycle constraint directly into the learning objective. Our method is purely unsupervised, requiring no ground-truth solutions or reinforcement rewards. To achieve this, we introduce three elements. First, we employ an equivariant and invariant representation that preserves permutation symmetry while removing nuisance variations such as translation and rotation, ensuring that the model operates directly on the intrinsic structure of the problem. Second, we use dropout as a source of controlled stochasticity within the network, which improves robustness without introducing iterative refinement. Third, we adopt snapshot ensembling along a single training trajectory, allowing multiple functionally distinct solutions to be obtained without additional training cost.
@@ 5.1 Equivariant Coordinate Feature Extraction
We introduce an equivariant coordinate feature extractor that maps a set of planar points to per-point features while respecting the symmetries of the Euclidean group. The construction is fully deterministic, permutation-equivariant, and (almost everywhere) invariant to global translations and rotations.
Problem Setting
@@
Intrinsic Coordinates and Harmonic Features
@@
Equivariance and Invariance Properties
@@
$f_{\mathrm{GNN}}(T_{t}\cdot X,A(T_{t}\cdot X))=f_{\mathrm{GNN}}(X,A(X)).$ (20)
@@ 6 Training and Inference
We train our scattering-attention GNN using distributed data parallelism on multi-node GPU clusters. For TSP100 and TSP200, we launch runs on $2$ nodes with $4$ GPUs per node (8 GPUs total), using per-GPU batch size $256$ and the Adam optimizer with weight decay $2.5\times 10^{-5}$, dropout $0.1$, and a learning rate $8\times 10^{-3}$. We use 16-layer and 20-layer GNNs with temperature $\tau=3$ and hidden dimension of 256 for TSP 100 and TSP 200, respectively. We train for up to $1000$ epochs for TSP 100 and 200 and 500, with a scheduler and a $15$-epoch warmup, enable adaptive gradient clipping, and apply early stopping with patience $100$; checkpoints are saved every $5$ epochs with automatic resume enabled. For TSP 500, we scale to $4$ nodes with $4$ GPUs per node (16 GPUs total) and train a larger model (hidden dimension $384$, $48$ layers) with dropout $0.5$, learning rate $2\times 10^{-3}$, weight decay $10^{-4}$, and temperature $\tau=3.5$, using a smaller per-GPU batch size of $50$ due to memory constraints. Across all settings, training and validation instances are sampled from uniform Euclidean TSP distributions, and we fix the distance scale to $5.0$ to match the inference-time construction of the graph adjacency.
@@ 7 Results
Figures 2 show the tour-length distributions on the test set for 100-, 200-, and 500-node instances, using the model that achieves the lowest validation length across all hyperparameter settings. Across all problem sizes, our method consistently outperforms the Greedy Nearest Neighbor (NN) baseline, which constructs tours by iteratively selecting the closest unvisited node. The resulting histograms show that our model produces shorter and more tightly concentrated tour-length distributions, with mean lengths of $\mu=8.18$, $11.65$, and $18.66$, compared to Greedy NN’s $\mu=9.67$, $13.43$, and $20.76$, respectively.
@@
500 Mean length 18.66 18.70 18.71 18.70 18.71 18.71 18.71 18.70 18.71 18.70 18.52
Wins (%) 14.3 9.3 10.2 9.2 9.6 9.1 9.1 10.1 10.6 8.5
@@
500 Mean length 18.66 18.72 18.72 18.70 18.68 18.51
Wins (%) 29.8 15.4 16.3 17.1 21.4
@@
GNN, UL 8.11 4.65% 3.37ms 11.57 8.43% 9.48ms 18.52 12.45% 59.34ms
Ours (batch size = 512)
dropout + snapshot
GNN, UL 8.09 4.39% 4.70ms 11.55 8.25% 13.28ms 18.47 12.11% 82.63ms
Table 2 illustrates the behavior of our snapshot ensemble across problem sizes. Each column corresponds to a single inference run using the same model architecture, but taken from different training epochs. For TSP 100 and TSP 200, snapshots are drawn from training epochs $\{850,900,950,1000\}$, while for TSP 500 they are taken from $\{400,450,500,550\}$.
@@
Across all problem sizes, we observe a clear and smooth trade-off between solution quality and inference cost as different ensemble strategies are applied to the same unsupervised GNN. Even in the deterministic setting, a single forward pass already achieves strong performance, with optimality gaps of 5.55%, 9.18%, and 13.30% on TSP 100, TSP 200, and TSP 500, respectively, at sub-millisecond to few-millisecond latency. demonstrating that the model internalizes substantial global structure without search or iterative refinement. Snapshot ensembling over five checkpoints and MC-dropout ensembling over ten stochastic passes further and consistently reduce the gap by exploiting complementary sources of diversity arising from optimization dynamics and controlled representation-level stochasticity, respectively, with the combined strategy achieving the strongest results overall (4.39%, 8.25%, and 12.11%). Importantly, these gains are obtained without additional training runs, neighborhood exploration, or post hoc local improvement, and inference remains firmly in the millisecond regime. Compared to classical heuristics, this yields a consistently more favorable quality–efficiency trade-off: while Christofides attains gaps of 12.12%, 12.75%, and 13.30% on TSP 100, TSP 200, and TSP 500, respectively, our feed-forward GNN substantially outperforms it on smaller instances and matches it on TSP 500, at orders-of-magnitude lower inference cost on GPUs. More aggressive heuristics such as 2-opt can further reduce the gap, but only at the expense of significantly higher runtime, particularly as problem size grows. Overall, these results suggest that learned, structure-aware heuristics can rival or surpass classical hand-designed methods, while naturally inducing a continuous and controllable trade-off between solution quality and inference cost.